25. Find the midsection of a frustum of a square pyramid if its lower base is 20 cm on a side, its upper base is 12 cm on a side, and the altitude of the frustum is 14 cm.
1
Expert's answer
2020-12-17T07:17:20-0500

Given "a=20cm, b=12cm,h=14cm"


"\dfrac{x}{b/2}=\dfrac{x+h/2}{m/2}=\dfrac{x+h}{a/2}"

"mx=bx+bh\/2"

"ax=bx+bh"

"x=\dfrac{bh}{a-b}"

"m=b+\dfrac{bh}{2bh}(a-b)"

"m=\dfrac{20+12}{2}=16(cm)"

"A_m=m^2=(16)^2=256(cm^2)"

"256\ cm^2"